OpenStudy (anonymous):

In Session 17 at around 18:00 in lecture 6, how does the professor get \[\frac{ d }{dx }f'(kx) = k * f'(kx)\]? Was it supposed to be \[\frac{ d }{dx }f'(kx) = k * f'(x)\]? Please explain how and if this was a mistake also explain how.

OpenStudy (anonymous):

No, he is correct, but he uses the chain rule in a way you normally wouldn't think to use it. For example, lets differentiate \[y = 6x\] You know that differentiating any function in the form \[cx\] results in \[c\]. So \[\frac{ dy }{ dx } 6x = 6\] But lets also do this same problem using the chain rule. Let the inside function be \[u = x\] and so the outside function is \[y = 6u\] Using the chain rule, \[\frac{ dy }{ dx }y = \frac{ dy }{ du }\frac{ du }{ dx }6u \] We know \[\frac{ du }{ dx } = 1\] So now we have \[\frac{ dy }{ dx }y = \frac{ dy }{ du }6u \] so \[\frac{ dy }{ dx } = 6\] Which is the same result we got above. This is exactly what the professor is doing, using the chain rule on a constant. Sorry if my explanation wasn't completely clear, please ask if you still have any doubts because this is my first time answering!

OpenStudy (anonymous):

The key is to understand that the function is being applied to the constant. For example, in \[f(x)=2x^2\] the function doesn't do anything to the constant 2, so the derivative of this function is the same as 2 times the derivative of x squared. But if this is our function, what would we get if we applied it to 3x?\[f(3x)=2(3x)^2\]Now the function is being applied to the constant, and it would be a mistake to say the derivative of this function is the same as three times the derivative of 2 times x squared. That's why we need the chain rule in this situation.